Find an equation for the line with the given properties.
A) Parallel to the line y=-3x; containing the point (5,6)?
B) Perpendicular to the line y=4x+2; containing the point (2,2)?



Answer by lwsshak3(11628)

Find an equation for the line with the given properties.
**
A) Parallel to the line y=-3x; containing the point (5,6)?
Equation of a line: y=mx+b, m=slope, b=y-intercept
For given line y=-3x: slope=-3
Parallel line will have the same slope.
Equation: y=-3x+b
solve for b using coordinates of given point(5,6)
6=-3*5+b
b=21
Equation of parallel line: y=-3x+21
..
B) Perpendicular to the line y=4x+2; containing the point (2,2)?
For given line y=4x+2: slope=4
Perpendicular line line will have a slope=negative reciprocal of slope of given line=-1/4.
Equation: y=-x/4+b
solve for b using coordinates of given point(2,2)
2=-2/4+b
b=5/2
Equation of perpendicular line line: y=-x/4+5/2